OpenStudy (nicole):

how to solve logarithms: 7(exponent) x-3=25

OpenStudy (anonymous):

Please clarify: do you mean: \[7^{x-3}=25\] or: \[7^x-3=25\] Thank you.

OpenStudy (nicole):

The first

OpenStudy (anonymous):

Assuming you mean the top equation (the harder one), then: \[7^{x-3}=25 \implies \log{7^{x-3}} = \log{25} \implies (x-3)\log{7} = \log{25}\] \[ \implies x = [\log25/\log7] +3\]

OpenStudy (nicole):

Okay great, thank you!